compute fxy and fyx for the function f(x,y) =e^x+y sin x+9x^2+2xy at (1,2)
1
Expert's answer
2020-04-06T11:47:18-0400

Find first "f_x" and "f_y"

"f_x=\frac{\partial}{\partial x}(e^{x}+y\sin x+9x^2+2xy)=e^x+y\cos x+18x+2y," "f_y=\frac{\partial}{\partial y}(e^{x}+y\sin x+9x^2+2xy)=\sin x+2x."

To find "f_{xy}" and "f_{yx}", take the partial derivative of "f_x" with respect to "y" and "f_y" with respect to "x", respectively.

"f_{xy}=\frac{\partial}{\partial y}(e^{x}+y\cos x+18x+2y)=\cos x+2,"

"f_{yx}=\frac{\partial}{\partial x}(\sin x+2x)=\cos x+2."

Then,

"f_{xy}(1,2)=f_{yx}(1,2)=\cos 1+2."